# arXiv:2603.01649v1[quant-ph] 02 Mar 2026

### v1

@@ I Introduction
However, photon shaping techniques typically assume a resonant frequency matching between emitter and receiver nodes [24, 28]. Although this condition greatly simplifies the required time-dependent controls, it introduces an additional source of error in realistic implementations, as well as extra constraints in chip manufacturing. Moreover, the resonant assumption also limits the allowed primitive operations to route or exchange quantum information among potentially detuned remote nodes, limiting scalability. Hence, finding controls to shape frequency-tunable photons emerges as a key element to both enlarge the capabilities of quantum networks, and circumvent or even harness frequency detunings among nodes for addressability. In this regard, it is worth noting a recent work in Ref. [52], that has reported a successful generation of frequency-tunable photons to overcome frequency mismatches among remote chips employing broadband transfer resonators.
@@
In this work, we take a different route to generate frequency-tunable single photons as compared to [52]. Here we provide a closed-form and analytical expression of the control, derived from a microscopic model, that exactly shapes a sech-like photon detuned by an arbitrary amount with respect to its emitter frequency. This is achieved by reverse engineering the emitter dynamics of a simple three level system. Producing such frequency-tunable photons demands a time-dependent modulation of both amplitude and phase of the coupling, that is experimentally feasible [48, 37, 28, 30]. Note that our theory is model independent, without further assumptions beyond its standard form, i.e. for resonant photons, and therefore it does not require neither extra elements in the setup as in [52], compared to previously reported experiments [37, 28, 30], nor modified circuit parameters. Given that such three-level model is ubiquitous across many platforms, our derived controls to yield frequency-tunable photons might find applications in other setups. Although the controls are theoretically exact, their experimental realization is prevented in some cases. Indeed, we find that producing sech-like photons at maximum bandwidth for any non-zero detuning results in an exponentially increasing control, both in amplitude and phase. For reduced bandwidths, this singular behavior disappears, being the half maximum bandwidth best suited for its realization. We then build on this general theory to design protocols on a two-node superconducting circuit network. By leveraging the frequency tunability of the injected or absorbed photons, we show how to overcome potential frequency mismatches among nodes to implement a frequency-selective and deterministic quantum state transfer. In addition, we employ these controls for remote Bell state preparation either among distant and detuned nodes or within the same chip without the need to physically interact among them. Our numerical simulations, taking into account all the elements of the network and under realistic parameters, result in high fidelities, mainly limited by decoherence effects. The developed theoretical framework highlights the suitability of the designed controls to produce frequency-tunable photons with current technology, and the new possibilities that they offer for quantum networking.
@@ II Pulse control for detuned photon emission
This divergence disappears for any $\eta>1$. In those cases, the maximum amplitude is bounded, $\max_{t}|g(t,\delta,\eta>1)|=\sqrt{4\delta^{2}\eta^{2}+(\eta-1)^{2}\kappa^{2}}/(2\eta\sqrt{\eta-1})$, which for $|\delta|\gg\kappa$ simplifies to $\max_{t}|g(t,\delta\gg\kappa,\eta>1)|\approx|\delta|/\sqrt{\eta-1}$. At the same time, the phase acquires a smooth behavior. This can be seen for example in the rate of the phase. Indeed, from Eq. (7), it follows that $\left.\dot{\varphi}(t,\delta,\eta>1)\right|_{t\rightarrow-\infty}=\delta$, and similarly $\left.\dot{\varphi}(t,\delta,\eta>1)\right|_{t\rightarrow+\infty}=\delta(\eta-2)/(\eta-1)$. For intermediate times, $\infty<t<\infty$, the rate $\dot{\varphi}(t,\delta,\eta>1)$ changes smoothly between these values (see App. C for further details). In Fig. 2(c) and (d) we provide a representation of the control, amplitude and phase, respectively, corresponding to photons with detuning $\delta=-2\kappa$ and reduced bandwidths, $\eta=3/2$ and $\eta=2$. The maximum value of $|g(t,\delta,\eta)|$ is depicted in Fig. 2(c) with horizontal lines for both cases. Note the significant difference in magnitude of the control with respect to $\eta=1$ (cf. Fig. 2(a) and (b)). This is precisely the motivation to consider detuned photons with a reduced bandwidth. While they demand the ability to tune both amplitude and phase of a time-dependent driving, the magnitudes of $|g(t,\delta,\eta)|$ and $\dot{\varphi}(t,\delta,\eta)$ are similar to the targeted detuning frequency $\delta$. These requirements are already met in current technology (see for example [37, 48]). Therefore, this opens the door for its experimental realization, unlocking potential applications where bridging frequency detunings between different elements may be critical or beneficial, as we further elaborate in Sec. III for a waveguide-QED quantum network. However, before moving on to discuss quantum networking protocols based on detuned single photons, in Sec. II.1 we further analyze the spectral properties of the controls and how realistic imperfections may impact the quality of the generated photons.
@@ II.1 Photon-emission fidelity
Figure 3: (a) Photon-emission infidelity $1-F_{e}(\tau)$ for a finite emission time $\tau$ as a function of the rescaled time $\kappa\tau/\eta$. Solid (black) line corresponds to the theoretical expression $1-F_{e}(\tau)=1-\tanh^{2}(\kappa\tau/(4\eta))$, while points show the numerically-computed infidelity for different photon detuning $\delta$ and reduced bandwidth $\eta$. Panel (b) shows the impact on the photon-emission infidelity as a function of the maximum allowed driving amplitude $g_{m}$ rescaled by $\delta$, for a fixed time $\kappa\tau=40$, and $\eta=1.25$ (orange), $\eta=2$ (blue) and $\eta=3$ (green). Solid (dashed) lines correspond to $\delta=2\kappa$ ($\delta=4\kappa$). (c) Spectra $|G(\omega,\delta,\eta)|^{2}$ of the controls for $\delta=4\kappa$, rescaled to their maximum, for three cases $\eta=3/2$ (orange circles), 22 (blue diamonds) and 33 (green triangles). Note the low-frequency components for $\eta=2$, and central frequencies located at $\omega=\delta(\eta-2)/(\eta-1)$. (d) Colormap of the difference between filtered with $\omega_{co}=10\kappa$ and ideal controls quantified by $\log_{10}S$ (cf. Eq. (10)) as a function of $\delta$ and $\eta$. The colormap encodes the scale of $\log_{10}S$, i.e. $S$ ranging from $10^{-5}$ (blue) to 11 (red).
@@
As commented above, another important aspect to quantify is the existence of a maximum achievable driving amplitude, $g_{m}$. This may important to ensure either the validity of the RWA, or to meet experimental constraints. A fixed $g_{m}$ hinders the exact generation of detuned photons when $g_{m}<\max_{t}|g(t,\delta,\eta)|$. To study this effect, we simply truncate the exact control. That is, if at some time $t$ the exact modulus (cf. Eq. (6)) surpasses $g_{m}$, i.e. $|g(t,\delta,\eta)|>g_{m}$ , its amplitude is instead replaced by $g_{m}$, while leaving the phase unaltered. The numerical results are shown in Fig. 3(b), where $1-F_{e}(g_{m})$ is plotted as a function of $g_{m}/\delta$ for a fixed protocol time $\kappa\tau=40$. The sharp drop in infidelity takes place when $g_{m}\approx\max_{t}|g(t,\delta,\eta)|$. From that point on, increasing $g_{m}$ does not improve $F_{e}$, as it remains limited by the finite time $\tau$ (cf. Fig. 3(a)). Note that although smaller $\eta$ values demand shorter emission times, they do so at the expense of an increasingly stronger driving. Recall that the maximum required amplitude for an exact emission for $\eta\rightarrow 1$ grows according to $|\delta|/\sqrt{\eta-1}$. This trade-off may be relevant in setups where strong drivings are either discouraged or infeasible, but longer emission times are within reach, or vice versa. It is worth noting that $\eta=2$ emerges as a good balance between emission time and required control amplitude, necessitating $g_{m}\approx\delta$. Indeed, the suitability of the $\eta=2$ case is also grounded in the spectral properties of $g(t,\delta,\eta)$, which we analyze below.
@@ III.4 Remote Bell state preparation
Figure 6: Panels (a) and (b) show the controls and population dynamics for preparing a Bell state among $q_{0}$ and $q_{1}$. In panel (a) $g_{2}(t)=0$ (not shown), and the control $g_{0}(t)$ (green line) is multiplied by 1010 for a better visualization. The real and imaginary components of $g_{1}(t)$ (and of $g_{2}(t)$ in (c)) are plotted with solid and dashed blue (red) lines, respectively. Similarly, panels (c) and (d) show the controls and dynamics for a Bell state among $q_{1}$ and $q_{2}$, with the same format as in (a) and (b). Again, in (c) the control $g_{0}(t)$ is multiplied by $10$. In all cases, $\kappa=2\pi\times 30$ MHz, $\eta=2$, $\omega_{0}=2\pi\times 8.5$ GHz and $\omega_{1,2}=\omega_{0}\pm\delta$ with $\delta=2.5\kappa$. Vertical gray lines mark the central emission and absorption times for each case. See main text for further details.
@@ Acknowledgements.
Á.P. acknowledges support from the CSIC JAE-Intro programme (ref. JAEINT_25_03528). Á.G.-L. acknowledges support by MICIU/AEI/10.13039/501100011033 and by ERDF/EU to project PID2023-146531NA-I00 and from the QUANTERA project MOLAR with Project No. PCI2024-153449 funded by Grant No. MICIU/AEI/10.13039/501100011033. R.P. acknowledges financial support from the Spanish Government via the project PID2024-161371NB-C21 (MCIU/AEI/FEDER, EU), and the Ramón y Cajal (RYC2023-044095-I) research fellowship.

### v2

# Shaping frequency-tunable single photons for quantum networking in waveguide QED

@@ I Introduction
However, photon shaping techniques typically assume a resonant frequency matching between emitter and receiver nodes [24, 28]. Although this condition greatly simplifies the required time-dependent controls, it introduces an additional source of error in realistic implementations, as well as extra constraints in chip manufacturing. Moreover, the resonant assumption also limits the allowed primitive operations to route or exchange quantum information among potentially detuned remote nodes, limiting scalability. Hence, finding controls to shape frequency-tunable photons emerges as a key element to both enlarge the capabilities of quantum networks, and circumvent or even harness frequency detunings among nodes for addressability. In this regard, it is worth noting recent works in Refs. [52, 53], that has reported a successful generation of frequency-tunable photons to overcome frequency mismatches among remote chips employing broadband transfer resonators.
@@
In this work, we take a different route to generate frequency-tunable single photons as compared to [52, 53]. Here we provide a closed-form and analytical expression of the control, derived from a microscopic model, that exactly shapes a sech-like photon detuned by an arbitrary amount with respect to its emitter frequency. This is achieved by reverse engineering the emitter dynamics of a simple three level system. Producing such frequency-tunable photons demands a time-dependent modulation of both amplitude and phase of the coupling, that is experimentally feasible [48, 37, 28, 30]. Note that our theory is model independent, without further assumptions beyond its standard form, i.e. for resonant photons, and therefore it does not require neither extra elements in the setup as in [53], compared to previously reported experiments [37, 28, 30], nor modified circuit parameters. Given that such three-level model is ubiquitous across many platforms, our derived controls to yield frequency-tunable photons might find applications in other setups. Although the controls are theoretically exact, their experimental realization is prevented in some cases. Indeed, we find that producing sech-like photons at maximum bandwidth for any non-zero detuning results in an exponentially increasing control, both in amplitude and phase. For reduced bandwidths, this singular behavior disappears, being the half maximum bandwidth best suited for its realization. We then build on this general theory to design protocols on a two-node superconducting circuit network. By leveraging the frequency tunability of the injected or absorbed photons, we show how to overcome potential frequency mismatches among nodes to implement a frequency-selective and deterministic quantum state transfer. In addition, we employ these controls for remote Bell state preparation either among distant and detuned nodes or within the same chip without the need to physically interact among them. Our numerical simulations, taking into account all the elements of the network and under realistic parameters, result in high fidelities, mainly limited by decoherence effects. The developed theoretical framework highlights the suitability of the designed controls to produce frequency-tunable photons with current technology, and the new possibilities that they offer for quantum networking.
@@ II Pulse control for detuned photon emission
This divergence disappears for any $\eta>1$. In those cases, the maximum amplitude is bounded. The long-time amplitude becomes $|g(t\rightarrow\infty,\delta,\eta>1)|=\sqrt{4\delta^{2}\eta^{2}+(\eta-1)^{2}\kappa^{2}}/(2\eta\sqrt{\eta-1})$, which also corresponds to its maximum value when $|\delta|\gtrsim\kappa$, so that $\max_{t}|g(t,|\delta|\gtrsim\kappa,\eta>1)|\approx|\delta|/\sqrt{\eta-1}$. At the same time, the phase acquires a smooth behavior. This can be seen for example in the rate of the phase. Indeed, from Eq. (7), it follows that $\left.\dot{\varphi}(t,\delta,\eta>1)\right|_{t\rightarrow-\infty}=\delta$, and similarly $\left.\dot{\varphi}(t,\delta,\eta>1)\right|_{t\rightarrow+\infty}=\delta(\eta-2)/(\eta-1)$. For intermediate times, $\infty<t<\infty$, the rate $\dot{\varphi}(t,\delta,\eta>1)$ changes smoothly between these values (see App. C for further details). In Fig. 2(c) and (d) we provide a representation of the control, amplitude and phase, respectively, corresponding to photons with detuning $\delta=-2\kappa$ and reduced bandwidths, $\eta=3/2$ and $\eta=2$. The maximum value of $|g(t,\delta,\eta)|$ is depicted in Fig. 2(c) with horizontal lines for both cases. Note the significant difference in magnitude of the control with respect to $\eta=1$ (cf. Fig. 2(a) and (b)). This is precisely the motivation to consider detuned photons with a reduced bandwidth. While they demand the ability to tune both amplitude and phase of a time-dependent driving, the magnitudes of $|g(t,\delta,\eta)|$ and $\dot{\varphi}(t,\delta,\eta)$ are similar to the targeted detuning frequency $\delta$. These requirements are already met in current technology (see for example [37, 48]). Therefore, this opens the door for its experimental realization, unlocking potential applications where bridging frequency detunings between different elements may be critical or beneficial, as we further elaborate in Sec. III for a waveguide-QED quantum network. However, before moving on to discuss quantum networking protocols based on detuned single photons, in Sec. II.1 we further analyze the spectral properties of the controls and how realistic imperfections may impact the quality of the generated photons.
@@ II.1 Photon-emission fidelity
Figure 3: (a) Photon-emission infidelity $1-F_{e}(\tau)$ for a finite emission time $\tau$ as a function of the rescaled time $\kappa\tau/\eta$. Solid (black) line corresponds to the theoretical expression $1-F_{e}(\tau)=1-\tanh^{2}(\kappa\tau/(4\eta))$, while points show the numerically-computed infidelity for different photon detuning $\delta$ and reduced bandwidth $\eta$. Panel (b) shows the required maximum amplitude of the control $g_{m}$ to emit a photon, as function of the detuning $|\delta|$, with fidelities $F_{e}=0.9$, $0.99$ and $0.999$, corresponding to light to darker points, and for different reduced bandwidths $\eta=1.25$ (orange), $\eta=2$ (blue) and $\eta=3$ (green), with $\kappa\tau/\eta=30$. Solid lines show the maximum value for a perfect emission when $|\delta|\gtrsim\kappa$, i.e. $g_{m}=|\delta|/\sqrt{\eta-1}$. (c) Spectra $|G(\omega,\delta,\eta)|^{2}$ of the controls for $\delta=4\kappa$, rescaled to their maximum, for three cases $\eta=3/2$ (orange circles), 22 (blue diamonds) and 33 (green triangles). Note the low-frequency components for $\eta=2$, and central frequencies located at $\omega=\delta(\eta-2)/(\eta-1)$. (d) Colormap of the difference between filtered with $\omega_{co}=10\kappa$ and ideal controls quantified by $\log_{10}S$ (cf. Eq. (10)) as a function of $\delta$ and $\eta$. The colormap encodes the scale of $\log_{10}S$, i.e. $S$ ranging from $10^{-5}$ (blue) to 11 (red).
@@
As commented above, another important aspect to quantify is the existence of a maximum achievable driving amplitude, $g_{m}$. This may important to ensure either the validity of the RWA, or to meet experimental constraints. A fixed $g_{m}$ hinders the exact generation of detuned photons when $g_{m}<\max_{t}|g(t,\delta,\eta)|$. To quantify this effect, we compute the maximum value $g_{m}$ that ensures a targeted emission fidelity $F_{e}$, by simply truncating the exact control until the fidelity equals the targeted value, while leaving the phase unaltered. In addition, we ensure that finite-time truncation is negligible setting $\kappa\tau/\eta=30$. The numerical results are shown in Fig. 3(b), where we show $g_{m}/\kappa$ as a function of the photon detuning $|\delta|/\kappa$ for three different targeted fidelities $F_{e}=0.9$, $0.99$ and $0.999$, corresponding to light to darker points, and for three different reduced bandwidths $\eta=1.25$, $2$ and $3$. As mentioned before, the maximum amplitude to ensure a perfect emission when $|\delta|\gtrsim\kappa$ simplifies to $\max_{t}|g(t,\delta,\eta>1)|\approx|\delta|/\sqrt{\eta-1}$, which are shown as solid lines in Fig. 3(b). As expected, as the targeted fidelity approaches one, the numerically computed values for $g_{m}$ tend to $|\delta|/\sqrt{\eta-1}$. This reflects the trade-off between strong drivings and protocol time. The smaller $\eta$, the larger the photon frequency bandwidth, and the faster the emission at the expense of larger amplitudes $g_{m}$. Conversely, the larger $\eta$, the smaller the required amplitude $g_{m}$, yet requiring longer protocol times. This trade-off may be relevant in setups where strong drivings are either discouraged or infeasible, but longer emission times are within reach, or vice versa. It is worth noting that $\eta=2$ emerges as a good balance between emission time and required control amplitude, necessitating $g_{m}\approx|\delta|$. Indeed, the suitability of the $\eta=2$ case is also grounded in the spectral properties of $g(t,\delta,\eta)$, which we analyze below.
@@ III.4 Remote Bell state preparation
Figure 6: Panels (a) and (b) show the controls and population dynamics for preparing a Bell state among $q_{0}$ and $q_{1}$. In panel (a) $g_{2}(t)=0$ (not shown), and the control $g_{0}(t)$ (green line) is multiplied by $10$ for a better visualization. The real and imaginary components of $g_{1}(t)$ (and of $g_{2}(t)$ in (c)) are plotted with solid and dashed blue (red) lines, respectively. Similarly, panels (c) and (d) show the controls and dynamics for a Bell state among $q_{1}$ and $q_{2}$, with the same format as in (a) and (b). Again, in (c) the control $g_{0}(t)$ is multiplied by $10$. In all cases, $\kappa=2\pi\times 30$ MHz, $\eta=2$, $\omega_{0}=2\pi\times 8.5$ GHz and $\omega_{1,2}=\omega_{0}\pm\delta$ with $\delta=2.5\kappa$. Vertical gray lines mark the central emission and absorption times for each case. See main text for further details.
@@ Acknowledgements.
The authors are grateful to Alonso Hernández-Antón for useful discussions and comments on the article. Á.P. acknowledges support from the CSIC JAE-Intro programme (ref. JAEINT_25_03528). R.P. acknowledges financial support from the Spanish Government via the project PID2024-161371NB-C21 (MCIU/AEI/FEDER, EU), and the Ramón y Cajal (RYC2023-044095-I) research fellowship.